3x-2y=3 ;2x+y=16 find dx and dy
Answer by Alan3354(69443)

3x-2y=3 ;2x+y=16 find dx and dy
-----------
Do you mean to differentiate implicitly?
3x-2y=3
3dx - 2dy = 0
---
2x+y=16
2dx + dy = 0